### v2

# Learning Hidden Physics and System Parameters with Deep Operator Networks

Abstract
Discovering hidden physical laws and identifying governing system parameters from sparse observations are central challenges in computational science and engineering. Existing data-driven methods, such as physics-informed neural networks (PINNs) and sparse regression, are limited by their need for extensive retraining, sensitivity to noise, or inability to generalize across families of partial differential equations (PDEs). In this work, we introduce two complementary frameworks based on deep operator networks (DeepONet) to address these limitations. The first, termed the Deep Hidden Physics Operator (DHPO), extends hidden-physics modeling into the operator-learning paradigm, enabling the discovery of unknown PDE terms across diverse equation families by identifying the mapping of unknown physical operators. The second is a parameter identification framework that combines pretrained DeepONet with physics-informed inverse modeling to infer system parameters directly from sparse sensor data. We demonstrate the effectiveness of these approaches on benchmark problems, including the Reaction-Diffusion system, Burgers’ equation, the 2D Heat equation, and 2D Helmholtz equation. Across all cases, the proposed methods achieve high accuracy, with relative solution errors on the order of $\mathcal{O}(10^{-2})$ and parameter estimation errors on the order of $\mathcal{O}(10^{-3})$, even under limited and noisy observations. By uniting operator learning with physics-informed modeling, this work offers a unified and data-efficient framework for physics discovery and parameter identification, paving the way for robust inverse modeling in complex dynamical systems.
keywords:
physics-deficient equation, deep operator network, system identification and generalization, scientific machine learning, inverse problem.
1 Introduction
Recent advances in machine learning, coupled with increasingly sophisticated data acquisition technologies, are transforming how we model and understand complex physical systems. In domains such as fracture analysis in electromechanical systems Tarafder et al. [2023], fluid dynamics Perdikaris and Karniadakis [2016] in aerospace engineering, or predicting cardiovascular blood flow in medical research Kennedy et al. [2017], Gennisson et al. [2013], and failure in materials Chen and Gu [2021], classical first-principles models often struggle to capture nonlinear dynamics or account for incomplete knowledge of governing processes. Sparse or noisy observations further challenge conventional numerical and statistical methods. As a result, there has been a growing interest in integrating machine learning with applied mathematics to bridge these gaps and accelerate scientific discovery. The landscape of data-driven dynamical systems discovery is rich and multifaceted, encompassing diverse methodological innovations. Researchers have developed approaches such as Equation-free modeling Kevrekidis et al. [2004], deep learning techniques Raissi [2018], non-linear regression Motulsky and Ransnas [1987], empirical dynamic modeling Chang et al. [2017], automated dynamics inference Daniels and Nemenman [2015], symbolic regression Koza [1992], and Koopman analysis Brunton et al. [2021] to address the fundamental challenge of modeling complex systems with incomplete knowledge and limited observational data.
A critical frontier in this domain is discovering closed-form mathematical models of real-world systems described by partial differential equations (PDEs) and identifying crucial system parameters from sparse, spatio-temporally scattered data. Traditional approaches have struggled to bridge the gap between theoretical understanding and practical measurement. Several prominent approaches have emerged to address these challenges. Existing methods like Sparse Identification of Non-linear Dynamics Brunton et al. [2016] (SINDy) and Deep Hidden Physics Models Raissi [2018] (DHPM) have made significant strides, but they face inherent limitations. SINDy approximates system behavior using sparse linear combinations of predefined candidate terms, but the framework requires prior knowledge of system non-linearities and is sensitive to measurement noise as well as is typically designed for recovering explicit governing equations for a single dynamical system. Deep Hidden Physics Models (DHPM) offers a promising alternative by utilizing two deep neural networks. One network captures the system state, while another approximates the unknown physics of the model. The DHPM framework demonstrates remarkable stability, leveraging automatic differentiation techniques to compute gradient terms, thus effectively bypassing the numerical differentiation challenges inherent in traditional approaches. DHPM offers improved stability through dual neural networks - one capturing system state, another approximating unknown physics - but cannot identify unknown system parameters from labeled datasets. Physics-Informed Neural Networks (PINNs) Raissi et al. [2019], Cai et al. [2021], Wang et al. [2024] marked a breakthrough in parameter estimation by introducing innovative functional approximation techniques. These methods simultaneously minimize both the residual loss from governing PDEs and the data loss from sparse sensor recordings, demonstrating success across diverse applications - from estimating hemodynamic parameters in cardiovascular systems Kissas et al. [2020] to reconstructing elasticity fields in heterogeneous materials Chen and Gu [2021], as well as inferring flow parameters in fluid mechanics Raissi et al. [2020], Zhang et al. [2023]. However, a fundamental limitation persists: existing methods typically require re-training and re-evaluation for each system parameter variation, limiting their generalizability and computational efficiency.
These challenges motivate the development of more flexible frameworks that (i) generalize across families of PDEs, (ii) infer hidden physics and system parameters simultaneously, and (iii) remain robust under sparse and noisy observations. Neural operators Lu et al. [2021], Goswami et al. [2022], Li et al. [2021], Tripura and Chakraborty [2023], Raonic et al. [2023], Cao et al. [2024] offer a compelling solution by learning mappings between infinite-dimensional functional spaces, thereby enabling operator-level generalization across different initial conditions, boundary conditions, or source terms. Among these, deep operator networks (DeepONet) Lu et al. [2021] are particularly appealing due to their architectural flexibility and proven success in modeling high-dimensional dynamical systems Kontolati et al. [2024], Agarwal et al. [2025], Wang et al. [2025], Sarkar et al. [2025b]. Recently, Molinaro et al. Molinaro et al. [2023] proposed a Neural Inverse Operator (NIO) methodology that learns operator-to-function mappings for canonical inverse problems, including Electrical Impedance Tomography, wave scattering, and seismic imaging. Their approach leverages multiple boundary measurements to characterize the observation operator and requires paired datasets of measurements and ground truth parameters generated from numerical simulations. Similarly, Long et al. Long et al. [2025] introduced an invertible Fourier Neural Operator (iFNO) that jointly learns forward and inverse mappings through a reversible architecture combined with variational autoencoders, enabling bidirectional predictions given complete paired training data. While these methods demonstrate the potential of neural operators for inverse problems, they fundamentally rely on the availability of labeled parameter datasets or complete field observations for training. In contrast, our work addresses scenarios where ground truth parameters are unavailable and only sparse sensor measurements are accessible, leveraging physics-informed constraints to enable parameter discovery directly from governing equations rather than paired supervision. In this study, we introduce two complementary physics-informed operator learning frameworks, built upon recent advances in neural operators:
  1. 1.
    The Deep Hidden Physics Operator (DHPO) network: The framework integrates DHPM with DeepONet to discover unknown governing physics directly from sparse data. Unlike DHPM, DHPO generalizes across multiple PDE families and system conditions, while maintaining interpretability by approximating the hidden terms of the governing equations.
  2. 2.
    Physics-Informed Parameter Identification: A framework for inferring system parameters (e.g., viscosity, diffusivity, thermal conductivity, source function) using sparse sensor recordings. Here, a pretrained DeepONet reconstructs the solution field from limited data, which then serves as input to a physics-constrained inverse model for parameter estimation. Unlike data-driven inverse operator approaches that require labeled parameter datasets, our framework addresses the more challenging scenario where ground truth parameters are unavailable during training. We leverage the known physics of the governing equations as constraints, enabling parameter identification purely from sparse sensor observations and PDE residuals without requiring supervised parameter labels.
By uniting operator learning with physics-informed modeling, this work contributes a step toward generalizable, data-efficient solvers for inverse problems in scientific computing. The proposed frameworks provide new opportunities for tackling hidden-physics discovery and parameter inference across diverse applications in engineering and the physical sciences. Both frameworks achieve high accuracy with limited observations, delivering relative $L_{2}$ errors on the order of $10^{-2}$ for hidden-physics discovery and absolute errors on the order of $10^{-3}$ for parameter identification. We demonstrate their performance on benchmark problems including the Burgers’ equation, a Reaction–Diffusion system, the 2D Heat equation, and the 2D Helmholtz equation and showing that our methods remain accurate and robust even under out-of-distribution scenarios. The remainder of the paper is organized as follows. Section 2 outlines the framework for building surrogates for hidden physics discovery and parameter identification. Section 3 demonstrates the applicability of the proposed framework using benchmark examples. Section 4 summarizes our study and highlights potential future research directions.
2 Physics Discovery and Parameter Identification using Neural Operators
The application of deep learning to the physical sciences hinges on two essential capabilities: uncovering governing equations directly from data when the underlying physics is only partially known, and estimating system parameters that control the dynamics of complex processes. Conventional data-driven approaches often fall short in these tasks, as they require large amounts of training data and generalize poorly across varying boundary conditions, source terms, or material properties. To address these limitations, we propose two complementary operator-learning frameworks based on DeepONet. The first, the Deep Hidden Physics Operator (DHPO), extends hidden-physics modeling into the operator-learning paradigm, enabling the discovery of unknown PDE terms across multiple equation families. The second is a parameter identification framework that leverages pretrained DeepONet to reconstruct solution fields from sparse sensor measurements and infer system parameters in a physics-consistent manner. Together, these approaches provide a unified, data-efficient strategy for physics discovery and parameter estimation in complex dynamical systems.
2.1 Deep Hidden Physics Operator (DHPO) - Discovering physics using DeepONet
Building upon the concept of deep hidden physics models (DHPM) introduced by Raissi et al. Raissi [2018], we introduce deep hidden physics operator (DHPO) - a framework that leverages the power of operator learning to discover underlying physical laws by identifying the mapping of unknown physical operators. While DHPM has proven effective in learning physics from data corresponding to a single PDE system under varying boundaries or initial conditions, our approach extends this capability to learn from multiple PDE systems simultaneously.
Consider a general nonlinear PDE of the form:
$\frac{\partial{u}}{\partial{t}}=\mathcal{N}(t,x,u,u_{x},u_{xx},..)+f(x),$ (1)
for $x\in\Omega,t\in[0,T],u:\Omega\times t\rightarrow\mathbb{R}$. Here, $\mathcal{N}$ represents the unknown physics operator we aim to discover, and $f(x)$ is a variable source term. Our framework, DHPO, employs a DeepONet architecture consisting of two primary components: a branch network that processes the input function $f(x)$ at fixed sensor locations, and a trunk network that handles the spatio-temporal coordinates $(x,t)$. The solution field $u(x)$ is approximated through the interaction of these networks via a dot product operation. The gradient, $u_{t},u_{x},u_{xx}$, is computed using the automatic differentiation approach. To discover the underlying physics, we introduce a hidden physics neural network that approximates $\mathcal{N}$ by taking the tuple $u,u_{x},u_{xx}]$ as inputs. This network learns to represent the unknown terms in the PDE while maintaining the interpretability of the discovered physics. It is important to note that interpretability here does not refer to symbolic equation recovery, but to mechanistic interpretability: identifying which physical processes are active, their relative contributions, and their functional dependence on solution gradients. The training of this framework is governed by a composite loss function that encompasses:
  1. 1.
    Initial condition constraints:
    $\mathcal{L}_{ic}=\frac{1}{N_{ic}b}\sum_{i=1}^{b}\sum_{j=1}^{N_{ic}}|u_{d}(x_{j}^{i},0)-u(x_{j}^{i},0)|^{2},$ (2)
  2. 2.
    Boundary condition constraints:
    $\mathcal{L}_{bc}=\frac{1}{N_{bc}b}\sum_{i=1}^{b}\sum_{j=1}^{N_{bc}}|u_{d}(0,t_{j}^{i})-u(0,t_{j}^{i})|^{2}+|u_{d}(1,t_{j}^{i})-u(1,t_{j}^{i})|^{2},$ (3)
  3. 3.
    PDE residual:
    $\mathcal{L}_{eqn}=\frac{1}{N_{Coll}b}\sum_{i=1}^{b}\sum_{j=1}^{N_{Coll}}|\left(\frac{\partial u}{\partial t}\right)_{(x_{j}^{i},t_{j}^{i})}-\left(\mathcal{N}(u,u_{x},u_{xx})\right)_{(x_{j}^{i},t_{j}^{i})}-f(x_{j}^{i})|^{2},$ (4)
  4. 4.
    Data fidelity:
    $\mathcal{L}_{data}=\frac{1}{N_{d}b}\sum_{i=1}^{b}\sum_{j=1}^{N_{d}}|u(x_{j}^{i},t_{j}^{i})-u_{d}(x_{j}^{i},t_{j}^{i})|^{2},$ (5)
where $u_{d}(x_{j}^{i},t_{j}^{i})$ and $u(x_{j}^{i},t_{j}^{i})$ denote the reference and predicted solutions at the $j^{th}$ point for the $i^{th}$ sample, respectively. The total loss function combines these components and is defined as:
$\mathcal{L}_{total}=\mathcal{L}_{ic}+\mathcal{L}_{bc}+\mathcal{L}_{eqn}+\mathcal{L}_{data}.$ (6)
The parameters of the trunk, branch, and hidden physics networks are optimized simultaneously by minimizing $\mathcal{L}_{total}$. A schematic of the DHPO framework is shown in Figure 1 and the detailed training procedure is presented in Algorithm 1.
Figure 1: Architecture of the proposed Deep Hidden Physics Operator framework developed to discover the unknown physics leveraging sparsely labeled dataset. The DeepONet learns the solution operator for varying system conditions to predict the desired output. The solution as well as the gradients are considered as inputs to the MLP and it predicts the unknown physics. The entire framework is trained in a single training session with physics loss as defined in Equation 4 and the data loss.
Our framework improves upon traditional DHPM through enhanced generalization that simultaneously handles multiple PDE systems across diverse input functions, boundary conditions, and geometries without retraining. It achieves improved data efficiency by leveraging physics-informed constraints and knowledge transfer, enabling learning from sparse or noisy measurements. The approach maintains increased interpretability, allowing direct extraction of governing equations and insights into physical term importance. Computational advantages emerge from physics-guided optimization, leading to faster convergence, better-conditioned loss landscapes, and reduced computational resources, delivering stable and consistent predictions across varying data scenarios and out-of-distribution conditions.
1Inputs: sensor measurements $f(x^{s})$ and $u(x^{s^{\prime}},t^{s^{\prime}})$ at $n$ fixed locations, collocation points $x,t$ , partially known physics $\mathcal{M}$, initial conditions, boundary conditions. $f(x)$ can be IC/BC/source or forcing $g(x)$.
2
3Initialize DeepONet $\mathcal{G}_{\bm{\theta}}$ with:
4     Branch network: process sensor inputs $f(x^{s})$
5     Trunk network: process coordinates $(x,t)$
6 Initialize MLP network for physics discovery with $\phi$
7 while not converged do
8 for each batch of sensor data do
9    Forward pass through DeepONet
10    $u_{\text{pred}}$ = $\mathcal{G}_{\bm{\theta}}(f(x^{s})(x,t))$
11    $u_{\text{ic}}$ = $\mathcal{G}_{\bm{\theta}}(f(x^{s})(x,0))$
12    $u_{\text{bc}}$ = $\mathcal{G}_{\bm{\theta}}(f(x^{s})(x,t))$
13    Compute gradients
14    $u_{x_{\text{pred}}}=du_{\text{pred}}/dx$
15    $u_{xx_{\text{pred}}}=d^{2}u_{\text{pred}}/dx^{2}$
16    $\cdots$
17    Forward pass through MLP
[…]
19    Compute loss
[…]
24    $\mathcal{L}(\bm{\theta},\phi)=\mathcal{L}^{\text{pde}}(\bm{\theta},\phi)(x,t)+\mathcal{L}^{\text{ic}}(\bm{\theta})(x,0)+\mathcal{L}^{\text{bc}}(\bm{\theta})(x,t)+\mathcal{L}^{\text{data}}(\bm{\theta})(x^{s^{\prime}},t^{s^{\prime}})$
25      Backpropagate and update $\bm{\theta}$ and $\phi$ using Adam
26   end for
27
28 end while
29Final optimized $\theta$, $\phi$
$\theta^{*}$, $\phi^{*}$ = argmin $\mathcal{L}(\theta,\phi)$
Algorithm 1 Training Scheme for the deep Hidden Physics Operator (DHPO).
We have considered the loss tolerance $\epsilon$ of $\mathcal{O}(10^{-3})$. To measure the deviation of the predicted solution from reference we use metric relative $L_{2}$ error and mean absolute error.
$\text{Relative}\ L_{2}\ \text{Error}=\frac{\sqrt{\sum_{i=1}^{N}|u_{i}-u_{i}^{*}|^{2}}}{\sqrt{\sum_{i=1}^{N}|u_{i}^{*}|^{2}}},$ (7)
$\text{Mean Absolute Error}=\frac{1}{N_{s}}{\sum_{i=1}^{N_{s}}E_{i};\text{where}\ E_{i}=\sum_{k=1}^{N_{i}}\frac{|u_{i}^{k}-u_{i}^{*k}|}{N_{i}}},$ (8)
where $u^{*},\;u$ are the reference and the predicted solution respectively. $N_{i}$ is number of data points for $i^{th}$ sample, $N_{s}$ is the total number of samples. We also define, the error distribution by $\mu\pm\sigma$, where $\mu=mean\{E_{i}\},\sigma=std.dev\{E_{i}\}$ are mean and standard deviation of errors over the samples. Unless otherwise noted, the mean test error refers to the mean absolute error (MAE) on the test set.
2.2 System Parameter Identification using DeepONet
2.2.1 Deterministic Parameter Estimation
Building upon our proposed architecture, we present a modified framework for parameter estimation in governing PDEs through surrogate modeling. While the underlying physics are known in this scenario, our objective is to determine the system parameters that govern the process (i.e., solving the parameter estimation inverse problem). Figure 2 illustrates the schematic of this modified architecture. For this framework, we have considered cases with just sensor recordings and no additional labeled datasets.
Figure 2: Schematic representation of the proposed architecture for unknown system parameters identification, $\nu$ for the Burger’s equation. In the schematic, $u^{s}$ denotes the known velocity field at fixed sensor locations. In step 1, a data-driven DeepONet is trained with loss defined at just the sensor locations to reconstruct the field based on sparse measurements. In the second step, the trained DeepONet is given a warm start, and the learning of the solution is improved in conjunction with an MLP which predicts the unknown $\nu$ employing a physics loss. $\bm{\theta}^{*},\bm{\phi}^{*}$ are the optimized parameters of the networks.
The framework consists of two key components. First, a DeepONet is pre-trained to learn the forward mapping from sparse input data to the complete solution field. Second, we introduce an additional MLP designed to learn the inverse mapping from the solution space to the parameter space. These two networks - the DeepONet and MLP - are then trained through continual learning, using the PDE residual as the primary loss term. To ensure solution accuracy and physical consistency, we incorporate additional loss terms for known data.
The fundamental advantage of this framework lies in its ability to simultaneously learn both forward and inverse mappings for any given physical system, without any additional labeled data beyond the sensor measurements. This dual-surrogate approach enables efficient parameter estimation while maintaining physical consistency through the embedded governing equations.
2.2.2 Probabilistic Parameter Estimation
The deterministic framework provides a single point estimate of the unknown parameter $\xi$. However, inverse problems often suffer from non-uniqueness where multiple parameter values can produce similar observations given sparse and noisy sensor data. To capture this inherent ambiguity, we extend our framework to estimate the full probability distribution of parameters. The MLP now outputs both the mean $\bm{\mu}_{\xi}$ and standard deviation $\bm{\sigma}_{\xi}$ of a Gaussian distribution: $\xi\sim\mathcal{N}(\bm{\mu}_{\xi},\text{diag}(\bm{\sigma}_{\xi}^{2}))$, where $\bm{\mu}_{\xi},\bm{\sigma}_{\xi}]=\mathcal{I}_{\phi}(\mathbf{u}^{s})$. During training, we employ the reparameterization trick ($\xi=\bm{\mu}_{\xi}+\bm{\sigma}_{\xi}\odot\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$) to enable backpropagation through the stochastic sampling process. The total loss function includes an additional KL divergence regularization term: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{pde}}(\bm{\theta},\bm{\phi})+\lambda_{\text{data}}\mathcal{L}_{\text{data}}(\bm{\theta})+\lambda_{\text{KL}}D_{\text{KL}}\left(\mathcal{N}(\bm{\mu}_{\xi},\text{diag}(\bm{\sigma}_{\xi}^{2}))\,\|\,p_{\text{prior}}(\xi)\right)$, where the KL term constrains the learned distribution to remain within physically meaningful bounds based on a prior $p_{\text{prior}}(\xi)$ informed by domain knowledge.
The probabilistic framework offers distinct advantages, it naturally quantifies epistemic uncertainty (narrow distributions indicate high confidence, broad distributions signal ambiguity), explicitly captures parameter non-uniqueness rather than forcing potentially misleading point estimates, accommodates measurement noise through learned variance that increases with data corruption, and enables risk-aware decision making through access to the full parameter distribution for uncertainty propagation. The training procedure follows Algorithm 2 with minor modifications: the MLP output dimension is doubled to predict both $\bm{\mu}_{\xi}$ and $\bm{\sigma}_{\xi}$, parameter sampling via reparameterization is performed before physics loss evaluation, and the KL divergence term is added to the total loss. In Section 3, we demonstrate both deterministic and probabilistic parameter identification, showing that the probabilistic framework provides calibrated uncertainty estimates that accurately reflect parameter ambiguity in sparse observation scenarios.
1Inputs and Training Data: sensor measurements $u(x^{s},t^{s})$ at $n$ fixed locations, collocation points $(x,t)$, PDE $\mathcal{N}(\mathcal{X},u(x,t),\xi)$, and training pairs $(u_{i}(x^{s},t^{s}),\xi_{i})_{i=1}^{N_{\text{train}}}$ consisting of $N_{\text{train}}$ system realizations with corresponding true parameters.
2 Step 1: Training solution operator $\mathcal{G}_{\bm{\theta}}$
3 Initialize DeepONet $\mathcal{G}_{\bm{\theta}}$ with:
4     Branch network: process sensor inputs $u(x^{s},t^{s})$
5     Trunk network: process coordinates $(x,t)$
6 while not converged do
7 for for each batch of K system realizations do
8    Forward pass through DeepONet at sensor locations
9      prediction = $\mathcal{G}_{\bm{\theta}}(u(x^{s},t^{s})(x^{s},t^{s}))$
10     Compute loss
11     $\mathcal{L}(\theta)=\frac{1}{K}\sum_{i=1}^{K}\left\lVert\text{prediction}_{i}-u_{i}(x^{s},t^{s})\right\rVert^{2}$
12
13     Backpropagate and update $\theta$ using Adam
14   end for
15
16 end while
17Store optimised $\bm{\theta}$
18     $\bm{\theta}^{*}=\text{ argmin}\mathcal{L}(\bm{\theta})$
19 Step 2: Learn parameter operator $\mathcal{I}_{\bm{\phi}}$
20 Initialize DeepONet $\mathcal{G}_{\bm{\theta}}$ with $\bm{\theta}=\bm{\theta}^{*}$ from step 1.
21 Initialize Parameter network $\mathcal{I}_{\phi}$: maps sensor data $\{u(x^{s},t^{s})\}$ to unknown parameter $\xi$
22 while not converged do
23 for each batch of data do
24    Forward pass through both networks
25    $u_{\text{pred}}$ = $\mathcal{G}_{\bm{\theta}}(u(x^{s},t^{s})(x,t))$
26    $\xi_{\text{pred}}=\mathcal{I}_{\bm{\phi}}(u(x^{s},t^{s}))$
27    Compute combine loss
[…]
30    $\mathcal{L}(\bm{\theta},\bm{\phi})=\mathcal{L}^{\text{pde}}(\bm{\theta},\bm{\phi})(x,t)+\mathcal{L}^{\text{data}}(\bm{\theta})(x^{s},t^{s})$
31      Back-propagate and update $\bm{\theta}$ and $\bm{\phi}$ using Adam
32   end for
33
34 end while
35
36Final optimized $\bm{\theta}$, $\bm{\phi}$
$\bm{\theta}^{*}$, $\bm{\phi}^{*}$ = argmin $\mathcal{L}(\bm{\theta},\bm{\phi})$
Algorithm 2 Training Scheme to obtain an operator framework for system parameter identification.
Algorithm 2 presents our two-stage training procedure for learning the parameter identification operator. The algorithm operates on a training dataset comprising multiple system realizations, where each realization corresponds to distinct parameter values and yields sensor measurements at fixed spatial locations over time. During training, the networks process mini-batches of these realizations, optimizing the loss functions through standard stochastic gradient descent until convergence.
3 Illustrative Problems
To examine our approach, we perform tests on four sets of examples including the Reaction Diffusion equation, Burger’s equation, Heat equation and Helmholtz Equation. The network architectures were systematically selected following established DeepONet guidelines Lu et al. [2021], with branch and trunk networks employing 3-4 hidden layers of 64-128 neurons. Hyperparameter sweep analysis demonstrates robust performance across reasonable architectural variations. Complete details are provided in Appendix A (Table A.1). For each problem, we demonstrate the applicability of our frameworks for hidden physics discovery, parameter identification, or both depending on the problem configuration.
3.1 Reaction Diffusion System
We first examine our approaches using a reaction-diffusion equation, which models various physical phenomena in nature such as oscillating chemical concentrations and spatial waves as seen in the Belousov-Zhabotinsky reaction, population dynamics, pattern formation in biological systems and morphology in crystals and alloys Smoller [2012]. The following equation describes the spatiotemporal evolution of the system and is given by:
[…]
where $D$ and $K$ denote the diffusion and the reaction coefficients, respectively. In Equation 9, IC denotes the initial condition and BC denotes the Dirichlet boundary condition. Going forward, we will first present the results of discovering the unknown physics using DeepONet, followed by presenting the results of employing DeepONet to characterize the diffusion coefficient of the system using sparse measurements.
Discovering physics using DeepONet We re-write Equation 9 to denote the known and the unknown part of the PDE, as
$\frac{\partial u}{\partial t}=\mathcal{N}(u,u_{x},u_{xx})+f(x),$ (10)
we assume that we know the time derivative of the problem and the remaining part of the governing equation shown in Equation 9 is unknown as well as approximated by $\mathcal{N}$, employing the proposed deep hidden physics framework within DeepONet. Our objective is to learn a solution operator $\mathcal{G}_{\bm{\theta}}$ that maps the source term $f(x)$ to the solution field $u(x,t)$ using partially known physics and sparse labeled data. For this setup, we have used $D=0.01$ $K=0.01$. To generate the training and testing data, the system in Equation 9 was solved using the finite difference method, with a spatial resolution of $\delta x=0.01$ units and a time step of $\delta t=0.001$ units. The final solution was evaluated on a $101\times 101$ grid in the space-time domain. We investigate three different input function spaces for source term $f(x)$. We train the model only using sine basis functions and test for all three function spaces.
  • Sine Basis Functions:
    $f(x)=\sum_{k=1}^{N_{f}}A_{k}sin(\pi kx),$ (11)
    where $A_{k}$ are coefficients from a Gaussian normal distribution with mean 0 and variance 1 and the number of frequencies $N_{f}=5$.
  • Gaussian Random Fields (GRFs):
    $GRF(x)\sim{\mathcal{G}(0,k_{l}(x_{1},x_{2}))},$ (12)
[…]
  • Modified GRFs:
    $\textnormal{Modified}\ GRF(x)=\alpha(x-x^{2})GRF(x)$ (13)
    where $\alpha=8.0$ ensures similar scale as sine basis functions.
To train the framework, the collocation points in the differential equation, initial conditions, and boundary conditions are defined as $N_{Coll}$, $N_{ic}$, and $N_{bc}$, respectively. These points are randomly generated for each sample using Latin hypercube sampling McKay et al. [1979], creating synthetic domain data that can vary at each gradient descent step. The training process is shown in Algorithm 1. For this problem, we use $N_{Coll}=2000$, $N_{ic}=200$, and $N_{bc}=250$. To determine the optimal amount of known labeled dataset required, we experiment with data point sets $N_{d}=\{200,500\}$ and training sample sets $N_{train}=\{50,100,200,500\}$, with a fixed number of test samples, $N_{test}=1000$.
The branch net takes the source term discretized at 101 spatial sensor locations, while the trunk net takes the spatial and temporal locations while the hidden physics MLP acts as a surrogate model for the right-hand side (RHS) of Equation 10. It accomplishes this by taking the solution $u$ and its spatial gradients, $u,u_{x},u_{xx}]$ as inputs and producing the prediction $\mathcal{N}$. The optimal model was found using training size of $N_{train}=500,N_{d}=500$. The loss curve over the training process is shown in Figure 1(a). The test mean absolute error for each of the cases is presented in Table 1, The solution is predicted within training time domain unless specifically mentioned.
Table 1: Reaction Diffusion system: Mean and standard deviation of the test error for all the three different input function spaces for the source term, $f(x)$. The experiments were carried out for five independent runs.
Method Test Error (Mean $\pm$ Std)
Sine basis $0.00871\pm 0.00424$
GRF $0.03918\pm 0.02095$
Modified GRF $0.01653\pm 0.00856$
Using sine basis functions for the source term yields the highest prediction accuracy among tested methods. Figure 4(a) illustrates the mean test error over various sample sizes and training data points, while Figure 4(b) shows the distribution of test errors for this optimal model with sine basis input function. We further evaluate the model’s predictive capability by using the learned hidden physics MLP $\mathcal{N}(x,t_{n};f)$ to solve the governing PDE via time integration beyond the training data (temporal extrapolation). Starting from the initial condition $u_{0}$, we apply forward Euler integration with the discovered physics term. For the Reaction-Diffusion system, the update rule is $u_{n+1}(x)=u_{n}(x)+(\mathcal{N}(x,t_{n};f)+f(x))\cdot dt$, while for Burgers’ equation it simplifies to $u_{n+1}(x)=u_{n}(x)+\mathcal{N}(x,t_{n};f)\cdot dt$. We compare two solution approaches: the direct DHPO prediction $u_{\text{pred}}$ (querying the trained DeepONet at future times) versus the time-integrated solution $u_{\text{hpm}}$ (obtained by stepping forward using the learned physics). We assess both approaches on unseen sine basis source terms as defined in Equation 11.
As shown in Figure 3, the model accurately predicts the DHPO solution ($u_{pred}$) and DHPO MLP integrated solution ($u_{hpm}$) at the training region. We also observe, in extrapolation region, the model accurately predicts the solution up to a specific time threshold, the accuracy in integrated solution is higher than DHPO solution, Hence we can infer that the Hidden Physics MLP is accurately capturing the dynamics. Figure 6 shows the contours of reference and predicted PDE terms. The inferred hidden physics term is given by,
$\mathcal{N}\approx D\frac{\partial^{2}u}{\partial x^{2}}+Ku^{2}.$
(a) Sample 1
(b) Sample 2
Figure 3: Reaction–diffusion equation: reference vs. DHPO predictions and HPM-integrated trajectories for two representative test samples drawn from the sine-function input space. Training data spans $t\leq 1$, while the extrapolation region ($t>1$) is indicated by white horizontal lines. For Sample 1, mean absolute errors are $u_{pred}(t\leq 1)=0.00432$, $u_{pred}(t>1)=0.04360$, $u_{hpm}(t\leq 1)=0.00226$, $u_{hpm}(t>1)=0.01917$. For Sample 2, mean absolute errors are $u_{pred}(t\leq 1)=0.01299$, $u_{pred}(t>1)=0.12473$, $u_{hpm}(t\leq 1)=0.00655$, $u_{hpm}(t>1)=0.05840$.
Unlike the sine basis, Gaussian random fields (GRFs) do not approach zero at the boundaries, leading to higher deviations in the predicted solution and physics terms, particularly at the boundaries (see Figure 4). Figure 5(a) illustrates the test error distribution for GRFs evaluated with the model trained with sine basis. In contrast, modified GRFs, which are zero at the boundaries similar to sine basis functions, achieve comparable prediction accuracy. Figure 5 shows that, with modified GRFs, the model accurately predicts both $u$ and the hidden physics terms, closely matching the ground truth. The test error distribution for modified GRFs is provided in Figure 5(b).
Additionally, we assess the model’s performance on modified GRFs by varying the length-scale parameter $l$ of the RBF kernel with values $l\in\{0.1,0.15,0.20,0.40,0.60\}$. As expected, Table 2 demonstrates that prediction error increases as the length scale decreases. However, for out-of-distribution functions with boundary values approaching zero, the model maintains relatively higher accuracy compared to cases with non-zero boundary conditions.
  • Noise robustness:
    Considering the impact of noise on model performance, an assessment was conducted using training data corrupted with synthetically generated random uniform noise. The noise was introduced as follows $u_{noise}=u_{clean}+U(0,\gamma\sigma_{u})$ , where $U$ is sampled from uniform random distribution with minimum value 0, maximum $\gamma\sigma_{u}$ , $\sigma_{u}$ denotes standard deviation in clean data, the parameter $\gamma$ controls the noise intensity. We train the model considering $\gamma\in\{0,0.2,0.5\}$, for direct comparison, both the training data volume and the training methodology were kept identical to those utilized for the optimal model presented earlier. The results are computed considering 1000 test inputs from sine basis function space. As illustrated in Table 3, the test error predictably increases with rising noise intensity. Notably, for noise intensity $\gamma<0.2$, the error exhibits only a marginal increase, $\approx 10\%$ . This indicates that the model maintains robustness in capturing the underlying dynamics within this noise threshold.
  • Sparsity:
    To evaluate the model’s performance under conditions of data sparsity, training was conducted using only two extreme temporal data points (t=0and t=1)(t=0\ $\text${and }t=1) along with 101 spatial data points. We train the model with 500 samples using the similar training procedure as shown in the previously trained optimal model. Evaluating the model on 1000 test samples yielded a mean test error of $0.00681\pm 0.00337$. When compared against the optimal model which was trained on 500 spatio-temporal data points, these results demonstrate that the sparse data model is remarkably capable of accurately predicting the dynamics, despite its limited training data.
    To assess the benefits of physics-informed training, we compare against a standard DeepONet trained on the same sparse data using identical branch and trunk network architectures. Evaluating on 1000 test samples, the DeepONet achieves a prediction error of $0.02619\pm 0.01074$ at the training time instants but degrades to $0.43818\pm 0.21707$ at interior (unobserved) time instants. This indicates severe overfitting to the training timestamps with poor generalization to the interior time domain. Representative predictions are shown in Figure 10.
Figure 4: Performance analysis for Reaction diffusion equation: (a) Mean test error over varying $N_{\text{train}}$. (b) Test error distribution of optimal model that achieved an error $0.00871\pm 0.00424$ with $N_{train}=500$, $N_{d}=500$.
Table 2: Reaction Diffusion system: Mean and standard deviation of the test error for all the case with input functions drawn from modified GRFs with varying length-scale parameter, $l$. The experiments were carried out for five independent runs.
$l$ Test Error
0.10 0.07945 $\pm$ 0.03666
0.15 0.03496 $\pm$ 0.01831
0.20 0.01653 $\pm$ 0.00856
0.40 0.00755 $\pm$ 0.00421
0.60 0.00656 $\pm$ 0.00378
Table 3: Reaction Diffusion system: Mean and standard deviation of the test error over the noise level.
noise intensity $\gamma$ Test Error
0.0 0.00871 $\pm$ 0.00424
0.2 0.00983 $\pm$ 0.00515
0.5 0.01652 $\pm$ 0.00895
Figure 5: Performance analysis for GRFs and Modified GRFs input function spaces: (a) GRFs basis $l=0.2$, Test error = 0.03918 $\pm$ 0.02095. (b) Modified GRFs basis $l=0.2$, Test error = 0.01653 $\pm$ 0.00856
System parameter identification: Diffusion coefficient
Deterministic. The neural operator framework introduced in Section 2.2 is employed to characterize the system, specifically to identify the diffusion coefficient of the governing PDE using solution values at selected spatial and temporal locations.
For this investigation, we considered varying diffusion coefficients and source terms in the Reaction-Diffusion equation defined in Equation 9. The equation was solved numerically using a finite difference method on a uniform grid with spatial resolution $\delta x=0.01$ and time step $\delta t=0.001$, creating a $101\times 101$ grid over the $x$-$t$ domain. The solution dataset was generated using 500 different diffusion coefficient values ($D$) uniformly distributed between $0.01$ and $0.05$. Each diffusion coefficient was paired with 20 distinct source terms $f(x)$ sampled from a Gaussian Random Field (GRF), resulting in 10,000 solution fields. The dataset was split into $N_{\text{train}}=8,500$ samples for training and $N_{\text{test}}=1,500$ samples for testing. To emulate sparse measurements, 300 spatial-temporal locations were randomly sampled and fixed across all training and testing samples, providing known values of $u(x,t)$. The network training hyperparameters are presented in Table 1. The training configuration employs $N_{\text{coll}}=2500$ collocation points for PDE residual evaluation. In the loss function, a penalty coefficient of $\lambda_{\text{coll}}=10$ was applied to the PDE residual term, while unity weights were maintained for data loss terms.
Figure 6 illustrates the framework’s capability in reconstructing the solution field $u(x,t)$ and estimating the diffusion coefficient $D$ from 300 sparse measurements. Results from two representative test cases demonstrate solution reconstruction with $L_{2}$ errors of order $\mathcal{O}(10^{-2})$, along with accurate recovery of the true diffusion coefficient values. Figure 7 presents the distribution of absolute errors in predicted diffusion coefficients across the entire test set, showing that the majority of predictions maintain absolute errors below $0.005$. This indicates the framework’s robust performance in parameter estimation tasks.
It is worth noting that determining the reaction coefficient from this system poses significant challenges due to the nonlinear relationship between $K$ and $u(x,t)$ in the reaction-diffusion equation. Small perturbations in the solution field $u(x,t)$ can lead to substantial errors in the estimation of $K$. Furthermore, the reaction term becomes negligible in regions where $u$ approaches zero, making the inverse problem for $K$ particularly challenging in these regions.
(a) Sample 1: Relative $L_{2}$ error of $u(x,t)$ = $0.027$. For this case, $D_{\text{true}}=0.027$ while $D_{\text{predicted}}=0.027$.
(b) Sample 2: Relative $L_{2}$ error of $u(x,t)$ = $0.029$. For this case, $D_{\text{true}}=0.036$ while $D_{\text{predicted}}=0.034$.
Figure 6: Reaction Diffusion Equation: Comparison of the reference solution and the predicted solution. The black cross marks represent the location where the data of $u(x,t)$ was known during the testing. The error at these specific locations is also shown in the absolute error plot.
Figure 7: Reaction Diffusion Equation: Distribution of absolute error of $D$ when compared to the $D$ used to generate the ground truth using the finite difference solver. The framework approximates $D$ with an error of $8\times 10^{-3}\pm 0.008$.
Probabilistic. To validate the probabilistic framework’s ability to capture parameter non-uniqueness, we evaluate whether parameter values sampled from the predicted distribution $\mathcal{N}(\mu_{D},\sigma_{D}^{2})$ can accurately reproduce the original sensor observations used as input to the inverse operator. For each test case, we sample 500 diffusion coefficient values from the learned distribution and perform forward simulations using a finite difference solver to obtain the corresponding solution fields. Figure 8 illustrates the predicted parameter distributions for two representative test samples, showing the mean estimate (dashed red line) alongside the true ground truth value (dashed green line). The distribution widths reflect the epistemic uncertainty inherent in the inverse problem given the available sparse sensor data. Figure 9 presents a scatter comparison between the true solution field $u$ (used to generate sensor measurements) and solutions obtained from sampled parameter values, demonstrating strong agreement across the spatial domain. Quantitatively, we achieve an average coefficient of determination $R^{2}=0.98$ across all sampled parameters, confirming that the inverse neural operator successfully captures the distribution of diffusion coefficients consistent with the observed data. This high reconstruction fidelity validates that the probabilistic framework appropriately quantifies parameter uncertainty while ensuring physical consistency.
Figure 8: Reaction Diffusion Equation: Predicted probability distributions of Diffusion coefficient $D$ for two representative test samples. The histogram represents 500 sampled values from the learned distribution $\mathcal{N}(\mu_{D},\sigma_{D}^{2})$. Dashed red line indicates the predicted mean, dashed green line shows the true Diffusion coefficient.
Figure 9: Reaction Diffusion Equation: Validation of predicted parameter distribution for representative samples through forward finite difference simulations. Each subplot shows the true solution field $u(x,t)$ (x-axis) versus the solution obtained using a Diffusion coefficient value sampled from the predicted distribution (y-axis).
3.2 Burgers’ equation
In this example, we consider the viscous Burgers’ equation defined as:
[…]
where $u(x,t)$ denotes the evolving spatio-temporal velocity field defined using the space and time coordinates, $(x,t)$. In Equation 14, BC denotes periodic boundary condition. Similar to the previous example, we will first present the results of discovering the unknown physics using DeepONet, followed by presenting the results of employing DeepONet to characterize the viscosity of the system using sparse measurements.
Discovering physics using DeepONet
In this problem, we re-write Equation 14 as:
$\frac{\partial u}{\partial t}(x,t)=\mathcal{N}(u,u_{x},u_{xx}),$ (15)
where the $\mathcal{N}$ denotes the hidden physics that is a function of the solution field $u(x,t)$ and its spatial derivatives $u_{x},u_{xx}$. In this problem, our aim is to learn the solution operator, $\mathcal{G}_{\theta}$ that maps the initial condition, $f(x,0)$ to the solution field, $u(x,t)$ using the partial known physics as well as sparse labeled dataset. In this case, we have considered fixed viscosity for the system, $\nu=0.01$. We solved equation (14) using a finite difference solver on a uniform grid ($\delta x=\delta t=0.01$), resulting in a $101\times 101$ grid spanning the $x$-$t$ domain with initial conditions sampled from a Gaussian Random Field (GRF).
The experiments were conducted with varying numbers of training samples, $N_{train}\in\{50,100,200,500,1000\}$ and the size of labeled dataset, $N_{d}\in\{200,500\}$, where $N_{d}$ denotes the number of locations per sample where the value of solution field is known. For testing, we used a fixed set of $N_{test}=1,000$ samples. Figure 10(a) illustrates the mean test error across different combinations of sample sizes and training data points. The optimal performance was achieved with $N_{train}=1000$ and $N_{d}=500$, resulting in mean absolute error of $0.00680\pm 0.00564$. The loss curve over the training process is shown in Figure 1(b).
Figure 10: Performance analysis for Burger’s equation: (a) Mean test error over varying $N_{\text{train}}$. (b) Test error distribution of optimal model that achieved an error $0.00680\pm 0.00564$ with $N_{train}=1000$, $N_{d}=500$.
The distribution of test errors for this optimal configuration is presented in Figure 10(b). As demonstrated in Figure 11, the model accurately predicts the solution when compared against ground truth. We also observe, in extrapolation region, the model accurately predicts the solution up to a specific time threshold, Hence we can infer that the Hidden Physics MLP is accurately capturing the dynamics, The Figure 9 shows the contours of PDE reference and predicted PDE terms, the model successfully approximated the unknown hidden term as,
$\mathcal{N}\approx 0.01\frac{\partial^{2}u}{\partial x^{2}}(x,t)-u(x,t)\frac{\partial u}{\partial x}(x,t)$
(a) Sample 1
(b) Sample 2
Figure 11: Burgers’ equation: reference vs. DHPO predictions and HPM-integrated trajectories for two representative test samples. Training data spans $t\leq 1$, and the extrapolation region ($t>1$) is indicated by white horizontal lines. For Sample 1, mean absolute errors are $u_{pred}(t\leq 1)=0.00250$, $u_{pred}(t>1)=0.00883$, $u_{hpm}(t\leq 1)=0.00456$, $u_{hpm}(t>1)=0.01198$. For Sample 2, mean absolute errors are $u_{pred}(t\leq 1)=0.01268$, $u_{pred}(t>1)=0.02325$, $u_{hpm}(t\leq 1)=0.01454$, $u_{hpm}(t>1)=0.02128$.
Relation to symbolic regression. For comparison, we applied weak-SINDy Messenger and Bortz [2021], a sparse symbolic regression method, to 300 randomly sampled measurements from a single trajectory, recovering $u_{t}=0.0101\cdot u_{xx}-1.0068\cdot u\cdot u_{x}$ with residual $3.95\times 10^{-3}$. While weak-SINDy efficiently discovers symbolic equations from limited data, DHPO learns generalizable operators of solution across varying initial conditions with physics discovery as an integrated component. The methods address complementary aspects of the inverse problem.
System parameter identification: Viscosity
Deterministic. We are now interested in employing the neural operator framework discussed in Section 2.2 to characterize the system, more specifically to identify the viscosity of the model while being provided with the solution values at a few spatial and temporal locations.
To design the problem, we have considered varying viscosity as well as initial conditions for the Burgers’ equation defined in Equation 14. For data generation, we solved this equation using a finite difference solver on a uniform grid with spatial resolution $\delta x=0.01$ and time step $\delta t=0.01$, resulting in a $101\times 101$ grid spanning the $x$-$t$ domain. The dataset comprises solutions corresponding to 500 different viscosity values ($\nu$) uniformly distributed between $0.01$ and $0.05$. Each viscosity value was paired with 20 distinct initial conditions sampled from a Gaussian Random Field (GRF), yielding a total of 10,000 solution fields. We randomly selected $N_{\text{train}}=8,500$ samples for training and reserved the remaining $N_{\text{test}}=1,500$ for testing the model. To simulate sparse measurements, we randomly sampled 300 spatial-temporal locations (fixed across all training and testing samples) from the solution field, where the value of $u(x^{s},t^{s})$ is known.
The hyperparameters involved in the training of the networks are presented in Table 1. The training configuration utilizes $N_{\text{coll}}=2500$ collocation points for evaluating the PDE residual. The loss functions were weighted with unity penalty coefficients.
Figure 12 demonstrates the framework’s performance in reconstructing both the solution field $u(x,t)$ and estimating the viscosity parameter $\nu$ from $300$ sparse measurements. We present results from two representative test cases selected from our test set of $1,500$ samples. The framework achieves solution reconstruction with $L_{2}$ errors of order $\mathcal{O}(10^{-2})$, while accurately recovering the true viscosity values in both cases. To provide a comprehensive assessment of the parameter estimation accuracy, Figure 13 presents the distribution of absolute errors in predicted viscosity values across the entire test set. The distribution reveals that the majority of predictions exhibit absolute errors below $0.002$, demonstrating the robust performance of our framework in parameter estimation tasks.
(a) Sample 1: Relative $L_{2}$ error of $u(x,t)$ = $0.092$. For this case, $\nu_{\text{true}}=0.028$ while $\nu_{\text{predicted}}=0.025$.
(b) Sample 2: Relative $L_{2}$ error of $u(x,t)$ = $0.061$. For this case, $\nu_{\text{true}}=0.032$ while $\nu_{\text{predicted}}=0.032$.
Figure 12: Burgers’ Equation: Comparison of the reference solution and the predicted solution. The black cross marks represent the location where the data of $u(x,t)$ was known during the testing. The error at these specific locations is also shown in the absolute error plot.
Figure 13: Burgers’ Equation: Distribution of absolute error of $\nu$ when compared to the $\nu$ used to generate the ground truth using the finite difference solver. The framework approximates $\nu$ with an error of $2\times 10^{-3}\pm 0.002$.
Probabilistic. We apply the same probabilistic framework to estimate the distribution of viscosity coefficient $\nu$ for Burgers’ equation. Following the validation procedure described above, we sample 500 viscosity values from the predicted distribution $\mathcal{N}(\mu_{\nu},\sigma_{\nu}^{2})$ for each test case and perform forward finite difference simulations to assess whether these sampled parameters can reproduce the original sensor observations. Figure 14 presents the predicted parameter distributions for two representative test samples. In both cases, the predicted mean (dashed red line) closely approximates the true viscosity value (dashed green line), while the distribution width reflects the epistemic uncertainty arising from sparse sensor measurements. Notably, Sample 2 exhibits a narrower distribution compared to Sample 1, indicating higher confidence in parameter estimation. This variation arises from differences in the information content of the respective sensor observations.
Figure 14: Burgers’ Equation: Predicted probability distributions of viscosity coefficient $\nu$ for two representative test samples. The histogram represents 500 sampled values from the learned distribution $\mathcal{N}(\mu_{\nu},\sigma_{\nu}^{2})$. Dashed red line indicates the predicted mean, dashed green line shows the true viscosity value.
Figure 15 validates the learned distributions through forward simulation: for each sample, we evaluate 5 different sampled viscosity values spanning the predicted distribution, and for each value, we compare the resulting solution field against the true field used to generate the input sensor data. The scatter plots demonstrate exceptional agreement across all sampled parameters, with coefficient of determination values exceeding $R^{2}>0.99$ in all cases confirms the physical consistency of the learned parameter distribution.
Figure 15: Burgers’ Equation: Validation of predicted parameter distribution for representative samples through forward finite difference simulations. Each subplot shows the true solution field $u(x,t)$ (x-axis) versus the solution obtained using a viscosity value sampled from the predicted distribution (y-axis).
To validate the posterior distribution, we perform forward finite difference simulations using sampled viscosity values and compare the resulting solutions against the reference solution obtained with the true viscosity. We distinguish between coefficient of determination ($R^{2}$) at the 300 sensor locations and relative $L_{2}$ error comparing the complete solution fields. Table 4 presents both metrics for 10 viscosity values sampled from the posterior for Sample 2.
The sampled values span a twofold range (0.0108 to 0.0212). When each sampled viscosity is used to solve Burgers’ equation forward, the resulting solutions exhibit full field $L_{2}$ errors from 0% to 17% relative to the true solution. This confirms that parameter variations significantly affect the complete PDE solution. However, at the 300 sensor locations, all sampled solutions achieve $R^{2}>0.95$ when compared to the reference sensor values, with most exceeding 0.99. Even the most extreme sample ($\nu=0.0108$, 44% below true value) attains $R^{2}=0.9571$ at sensors.
Table 4: Validation of posterior parameter distribution for Burgers’ equation (Sample 2). Each row shows a viscosity value sampled from $\mathcal{N}(\mu_{\nu},\sigma_{\nu}^{2})$. For each sample, we solve Burgers’ equation using finite differences and compare the resulting solution field against the reference solution obtained with the true viscosity ($\nu=0.01922$). The coefficient of determination measures agreement at the 300 sensor locations, while the relative $L_{2}$ error measures agreement over the complete $\times101$ spatio-temporal grid.
Sample Viscosity ($\nu$) $R^{2}$ at sensors Full field rel. $L_{2}$ error
True (ref) 0.01922 0.9990 0.0000
$\nu_{1}$ (mean) 0.01465 0.9863 0.0853
$\nu_{2}$ 0.01731 0.9961 0.0339
$\nu_{3}$ 0.01592 0.9919 0.0602
$\nu_{4}$ 0.01672 0.9945 0.0450
$\nu_{5}$ 0.02016 0.9995 0.0160
$\nu_{6}$ 0.01225 0.9705 0.1364
$\nu_{7}$ 0.01633 0.9933 0.0523
$\nu_{8}$ 0.02122 0.9992 0.0333
$\nu_{9}$ 0.01091 0.9583 0.1671
$\nu_{10}$ 0.01080 0.9571 0.1699
This behavior is the signature of ill-posed inverse problems with sparse observations. Multiple parameter configurations produce forward solutions that match observed sensor data equally well (high $R^{2}$) while differing substantially in the complete solution field (varying $L_{2}$ error). Solution variations concentrate in unobserved regions between sensors, where measurements provide no constraint. The posterior distribution correctly captures the range of viscosity values consistent with sparse observations, quantifying the epistemic uncertainty inherent to parameter estimation from limited data.
3.3 2D Heat equation
In the next example, we consider the heat equation over a L-shaped 2D computational domain $\Omega_{L}=\Omega\setminus R=\left[(-0.5,0.5)\times(-0.5,0.5)\right]\setminus\left[[0,0.5]\times[0,0.5]\right]$, defined as:
$\begin{split}&\frac{\partial u}{\partial t}(x,y,t)=\mathcal{K}(\frac{\partial^{2}u}{\partial x^{2}}(x,y,t)+\frac{\partial^{2}u}{\partial y^{2}}(x,y,t))\text{ on }\ \Omega_{L}\times[0,1],\\ \text{IC: }&u(x,y,0)=f(x,y),\\ \textnormal{BC: }&u(-0.5,y,t)=u(x<0,0.5,t)=u(0,y>0,t)=\\ &u(x>0,0,t)=u(0.5,y<0,t)=u(x,-0.5,t)=0\end{split}$ (16)
where $u(x,y,t)$ denotes the evolving spatio-temporal temperature field and $\mathcal{K}$ represents of the thermal conductivity coefficient. IC in Equation 16 denotes the initial condition sampled from GRF and BC represents the boundary condition. For data generation, we employed a finite element solver (deal.ii Arndt et al. [2021]) on a uniform grid with spatial resolution $\delta x\approx 0.03$ and time step $\delta t=0.01$, resulting in a total of 833 grid points covering the $x$-$y$ domain and 101 time steps. Similar to the previous examples, we will first present the results of discovering the unknown physics using DeepONet, followed by presenting the results of employing DeepONet to characterize the thermal conductivity coefficient of the system using sparse measurements.
Discovering physics using DeepONet
We learn the mapping between different ICs to the solution field using DHPO. We consider the RHS of Equation 16 unknown and re-write the equation as:
$\frac{\partial u}{\partial t}=\mathcal{N}(u,u_{x},u_{y},u_{xx},u_{yy}),$ (17)
where $\mathcal{N}$ is the hidden physics operator. The trunk network takes the spatio-temporal coordinates $(x,y,t)$ as input, while the branch network consists of a convolutional neural network (CNN) followed by a fully connected network (FCN). The CNN processes a $33\times 33$ image generated from the initial condition $f(x,y)$. The outputs of the trunk and branch networks are combined via a dot product, producing a scalar output that approximates the solution field $u$. In addition, we employ a hidden-physics neural network that receives the tuple $u,u_{x},u_{y},u_{xx},u_{yy}]$ as input and outputs a scalar quantity $\mathcal{N}$, representing the unknown PDE terms. Spatial derivatives are computed using reverse mode AD.
For training and validation, the thermal conductivity is fixed at $0.015$. A total of 500 training samples are generated by varying the initial condition. The overall loss function includes contributions from the initial condition loss, boundary condition loss, data loss, and equation residual loss. For the data loss, we sample $50/833$ spatial and $11/101$ uniformly spaced temporal points. The initial condition loss is evaluated at 200 fixed spatial locations across all samples. The boundary loss is computed at $400\times 100$ collocation points, corresponding to 400 spatial boundary locations at 100 time instants. The equation residual loss is evaluated at $500\times 100$ randomly sampled collocation points in space and time, with new samples drawn for each batch and epoch. The evolution of the total loss during training is shown in Figure 1(c). For testing, 1000 samples are evaluated, each on a grid of $833\times 101$ spatio-temporal points. The distribution of test errors is shown in Figure 16, yielding a mean absolute error of $0.01579\pm 0.00567$. Predictions for one representative test samples are presented in Figure 17, where the reconstructed solution fields closely match the reference, with mean absolute error of $0.01064$, respectively.
Figure 16: Heat equation: Test Error distribution using 1000 test samples, model achieved test mean absolute error of $0.01579\pm 0.00567$.
Figure 17: Heat equation: reference vs predicted solution field for a representative test sample (mean absolute error = 0.01064).
System parameter identification: Thermal conductivity coefficient
Deterministic. We now aim to apply the neural operator framework introduced in Section 2.2 to characterize the system by identifying the thermal conductivity coefficient, given access to solution values at a limited number of spatial and temporal locations.
To set up this problem, we consider a Heat equation model with varying thermal conductivity coefficients and initial conditions, as defined in Equation 16. The dataset includes solutions for 100 different thermal conductivity values ($\mathcal{K}$), uniformly sampled in the range $[0.01,0.2]$. Each value of $\mathcal{K}$ was paired with 100 distinct initial conditions sampled from a Gaussian Random Field (GRF), yielding a total of 10,000 solution instances. From these, we randomly selected $N_{\text{train}}=8,000$ samples for training and used the remaining $N_{\text{test}}=2,000$ samples for evaluation. To emulate sparse observations, we placed 40 sensors evenly across the spatial domain and recorded temperature values at 11 time steps within the interval $t\in[0,1]$.
The hyperparameters used during training are summarized in Table 1. The training setup includes $N_{\text{coll}}=2000$ collocation points for computing the PDE residual, and all loss terms are assigned unit penalty weights.
Figure 18: Heat’ Equation: Distribution of absolute error of $\mathcal{K}$ when compared to the $\mathcal{K}$ used to generate the ground truth using the finite element solver. The framework approximates $\mathcal{K}$ with an error of $3\times 10^{-2}\pm 0.04$.
Figure 19 illustrates the performance of our framework in reconstructing both the solution field $u(x,y,t)$ and the thermal conductivity coefficient $\mathcal{K}$ from a total of 440 sparse measurements. We display results from a representative test case drawn from the test set of 2,000 samples. The reconstruction achieves $L_{2}$ errors on the order of $\mathcal{O}(10^{-2})$, and the inferred values of $\mathcal{K}$ are reasonably accurate, despite the inverse problem’s inherent non-uniqueness. To further evaluate the accuracy of parameter estimation, Figure 18 shows the distribution of absolute errors in the predicted conductivity values across the test set. The framework approximates K with an error of $3\times 10^{-2}\pm 0.04$, confirming the robustness of our framework for thermal parameter identification. The high standard deviation relative to the mean reflects the inherent variability in inverse heat conduction problems, where parameter identifiability from sparse sensors depends strongly on the thermal conductivity magnitude and initial condition structure. The bimodal distribution shows that approximately $65\%$ of cases achieve errors below $0.02$, while more challenging cases with high conductivity values or low spatial variation in initial conditions exhibit reduced identifiability. The median error of $0.028$ provides a more robust characterization of typical performance, with 72% of predictions achieving errors below $0.03$.
Figure 19: Heat Equation: Comparison of reference solution and predicted solution. Black and white cross marks indicate sensor locations where u(x,y,t) data was available during testing. Sensor readings were collected at 11 time steps; no sensor data is shown at t=0.22, t=0.44 and t=0.66. The corresponding absolute errors at these sensor locations are displayed in the error plot.
Probabilistic. We apply the same probabilistic framework to estimate the distribution of thermal conductivity coefficient $K$ for the Heat equation. Following the validation procedure described above, we sample 500 thermal conductivity values from the predicted distribution $\mathcal{N}(\mu_{K},\sigma_{K}^{2})$ for each test case and perform forward finite element simulations to assess whether these sampled parameters can reproduce the original sensor observations. Figure 20 presents the predicted parameter distributions for two representative test samples. In both cases, the predicted mean (dashed red line) closely approximates the true thermal conductivity value (dashed green line), while the distribution width reflects the epistemic uncertainty arising from sparse sensor measurements distributed across the L-shaped domain. Both samples exhibit similar distribution widths ($\sigma_{K}\approx 0.002$), indicating comparable confidence levels in parameter estimation despite different initial temperature profiles.
Figure 21 validates the learned distributions through forward simulation: for each sample, we evaluate 10 different sampled thermal conductivity values spanning the predicted distribution, and for each value, we compare the resulting temperature field against the true field used to generate the input sensor data. The scatter plots demonstrate exceptional agreement across all sampled parameters, with coefficient of determination values reaching $R^{2}\geq 0.99$ in all cases, with several achieving perfect correlation ($R^{2}=1.0000$).
Figure 20: Heat Equation: Predicted probability distributions of thermal coefficient $\mathcal{K}$ for two representative test samples. The histogram represents 500 sampled values from the learned distribution $\mathcal{N}(\mu_{\mathcal{K}},\sigma_{\mathcal{K}}^{2})$. Dashed red line indicates the predicted mean, dashed green line shows the true thermal coefficient value.
Figure 21: Heat Equation: Validation of predicted parameter distribution for representative samples through forward finite difference simulations. Each subplot shows the true solution field $u(x,t)$ (x-axis) versus the solution obtained using a thermal coefficient value sampled from the predicted distribution (y-axis).
3.4 2D Helmholtz equation: Source reconstruction
In the next problem, we consider the steady state wave equation defined as:
[…]
where $u(x,y)$ denotes the steady-state wave field, $g(x,y)$ represents an unknown source distribution that we aim to reconstruct, and the wave number $k$ is set to $2\pi$. To make the inverse problem tractable, $g(x,y)$ is parameterized by a Gaussian function as:
$g(x,y)=A\exp\left(-\frac{(x-x_{0})^{2}+(y-y_{0})^{2}}{2\sigma^{2}}\right)$ (19)
where the standard deviation is fixed at $\sigma=0.2$, while the amplitude $A$ and center location $(x_{0},y_{0})$ are treated as unknown parameters to be determined from sparse measurements of $u(x,y)$. For data generation, we employed a Finite difference Helmholtz solver on a uniform grid with spatial resolution $\delta x=0.015$, resulting in $4096$ grid points. The dataset includes $800$ samples in which the three parameters $A,x_{0},\text{ and }y_{0}$ are sampled from uniform distributions over the bounds $[2,6]$, $[0.35,0.65]$, and $[0.35,0.65]$, respectively.
Deterministic. From the dataset, we randomly selected $N_{\text{train}}=640$ samples for training and used the remaining $N_{\text{test}}=160$ samples for testing. We placed $32\times 32$ uniformly spaced sensors across the domain to capture data for source reconstruction. The hyperparameters used during training are summarized in Table 1. The training setup includes $N_{\text{coll}}=4096$ collocation points for computing the PDE residual. We applied a penalty weight of $10$ to the data loss term, while all other loss terms were assigned unit penalty weights.
Figure 22 illustrates the distribution of absolute errors obtained in the recovery of the Helmholtz source parameters $A$, $x_{0}$, and $y_{0}$ across the entire test set. The inferred amplitudes exhibit a broader spread, with a mean absolute error of $0.28\pm 0.32$, reflecting the higher sensitivity of the solution field to variations in source strength. In contrast, the spatial parameters $x_{0}$ and $y_{0}$ are recovered with substantially smaller errors of $3.4\times 10^{-2}\pm 2.25\times 10^{-2}$ and $4.4\times 10^{-2}\pm 3.03\times 10^{-2}$, respectively. These results highlight the ability of the inverse model to accurately infer both the location and magnitude of the underlying Gaussian source from the observed PDE solutions, even in the presence of moderate variability in amplitude estimation.
Figure 22: Helmholtz Equation: Distribution of absolute errors for the inferred parameters $A$, $x_{0}$, and $y_{0}$. The mean absolute errors across 160 samples are $2.82\times 10^{-1}\pm 3.19\times 10^{-1}$ for $A$, $3.41\times 10^{-2}\pm 2.25\times 10^{-2}$ for $x_{0}$, and $4.44\times 10^{-2}\pm 3.03\times 10^{-2}$ for $y_{0}$.
The mean relative $L_{2}$ error for the predicted solution field $u(x,y)$ is $10.19\%$, while the source reconstruction achieves a mean relative $L_{2}$ error of $22.38\%$. For individual parameter estimation, the model achieves mean absolute errors of $0.33$ for amplitude $A$, $0.036$ for $x_{0}$, and $0.045$ for $y_{0}$, corresponding to mean relative errors of $6.85\%$, $7.68\%$, and $9.81\%$, respectively, indicating strong predictive performance for the amplitude and reasonable accuracy for the spatial location parameters. Figures 23 and 24 present comparisons of the reconstructed source and the corresponding solution prediction for a representative test sample.
Figure 23: Helmholtz equation: Source identification. The left panel shows the true Gaussian source $g(x,y)$, the middle panel shows the reconstructed source, and the right panel shows the pointwise absolute error. The reconstructed source attains a relative $L_{2}$ error of $0.2358$. For this sample, the ground-truth source parameters are: $A_{\text{true}}=2.4482$, $x_{\text{true}}=0.5237$, and $y_{\text{true}}=0.4031$. The predicted parameters are: $A_{\text{pred}}=2.4772$, $x_{\text{pred}}=0.5192$, and $y_{\text{pred}}=0.4699$. The corresponding absolute errors are: $|A_{\text{true}}-A_{\text{pred}}|=0.02896$, $|x_{\text{true}}-x_{\text{pred}}|=0.00451$, and $|y_{\text{true}}-y_{\text{pred}}|=0.06682$.
Figure 24: Helmholtz equation: Solution comparison. The left panel shows the FDM solution $u(x,y)$, the middle panel presents the predicted solution obtained from the learned operator, and the right panel displays the pointwise absolute error between the two fields. The solution field has a relative $L_{2}$ error of $0.1219$ and an MSE of $3.51\times 10^{-5}$.
Probabilistic. We apply the probabilistic framework to the Helmholtz source reconstruction problem, which involves estimating three parameters simultaneously: the amplitude $A$, and the center coordinates $(x_{0},y_{0})$ of a Gaussian source. This represents a more challenging multi-parameter inverse problem compared to the single-parameter cases above. Following the validation procedure, we sample 500 parameter sets from the predicted joint distribution, where each parameter follows $A\sim\mathcal{N}(\mu_{A},\sigma_{A}^{2})$, $x_{0}\sim\mathcal{N}(\mu_{x_{0}},\sigma_{x_{0}}^{2})$, and $y_{0}\sim\mathcal{N}(\mu_{y_{0}},\sigma_{y_{0}}^{2})$, and perform forward Finite difference Helmholtz simulations to assess data reconstruction capability.
Figure 25 presents the predicted distributions for all three parameters. The predicted means (dashed red lines) accurately capture the true parameter values (dashed green lines) for the spatial coordinates; however, the predicted mean for amplitude shows a noticeable offset from the true value. We hypothesize that this discrepancy might be due to a multimodal distribution of the amplitude for this problem. Notably, the spatial location parameters $x_{0}$ and $y_{0}$ exhibit relatively narrow distributions ($\sigma_{x_{0}}\approx 0.004$, $\sigma_{y_{0}}\approx 0.005$), indicating high confidence in source localization, while the amplitude parameter $A$ shows a broader distribution ($\sigma_{A}\approx 0.02$), reflecting greater uncertainty in estimating the source strength. This difference arises from the higher sensitivity of the wave field to source location compared to source amplitude, small spatial shifts produce more distinct measurement patterns than amplitude variations. Figure 26 validates the learned distributions through forward simulation: we sample 5 complete parameter sets $(A,x_{0},y_{0})$ from the joint distribution and compare the resulting wave fields against the true field. The scatter plots demonstrate exceptional agreement across all sampled parameter sets, with coefficient of determination values consistently exceeding $R^{2}>0.99$ for all cases, confirming that the probabilistic framework successfully captures the manifold of source parameters consistent with the observed wave field measurements.
Figure 25: Helmholtz Equation: Predicted probability distributions of Amplitude AA, x-coordinate of center $x_{0}$ and y-coordinate of center $y_{0}$ for two representative test samples. The histogram represents 500 sampled values from the learned distribution $\mathcal{N}(\mu_{A},\sigma_{A})$, $\mathcal{N}(\mu_{x_{0}},\sigma_{x_{0}})$ and $\mathcal{N}(\mu_{y_{0}},\sigma_{y_{0}})$. Dashed red line indicates the predicted means, dashed green line shows the true parameter values.
Figure 26: Helmholtz Equation: Validation of predicted parameter distribution for representative samples through forward finite difference simulations. Each subplot shows the true solution field $u(x,t)$ (x-axis) versus the solution obtained using a value sampled from the predicted distributions of parameters (y-axis).
Remark on parametric vs. non-parametric reconstruction. The parametric Gaussian representation in Equation 19 is a deliberate choice motivated by the fundamental ill-posedness of the inverse source problem. Direct reconstruction of the source field $g(x,y)$ on our $64\times 64$ computational grid would require recovering 4096 field values from only 1024 sensor measurements. This yields a severely underdetermined system with information ratio 1:4.
Such problems possess a nullspace of dimension at least 3072, meaning infinitely many distinct source distributions produce identical sensor observations Isakov [2006]. The singular value decomposition (SVD) of the forward operator reveals that high-frequency source components correspond to rapidly decaying singular values Hansen [2010]. These components generate evanescent waves that decay exponentially with distance, rendering them effectively invisible to sensors. Even modest measurement noise causes error amplification proportional to $1/\sigma_{\min}$, violating the discrete Picard condition Hansen [1987], Colton et al. [1998].
(a) MLP-based inverse operator: source reconstruction
(b) DeepONet-based inverse operator: source reconstruction
Figure 27: Helmholtz equation: Non-parametric source reconstruction comparison. (a) MLP-based approach with relative $L_{2}$ error of $0.3372$. (b) DeepONet-based approach with relative $L_{2}$ error of $0.5970$. For each subfigure: left panel shows the true Gaussian source $g(x,y)$, middle panel shows the reconstructed source, and right panel shows the pointwise absolute error. Both approaches correctly identify the approximate source location but fail to recover detailed spatial structure, confirming fundamental non-uniqueness.
In contrast, our parametric formulation reduces the problem to the well-posed space $\mathbb{R}^{3}$ with information ratio $341:1$. This exploits the physical prior that acoustic sources, antennas, and seismic events are typically spatially localized Bao and Li [2005], Li et al. [2017]. Alternative regularization strategies exist, including Tikhonov smoothness penalties Willoughby [1979], total variation constraints Rudin et al. [1992], and sparsity-based methods Candès et al. [2006]. However, all require comparable prior assumptions to make the problem well-posed. Our parametric representation offers the advantage of direct physical interpretability: the parameters $(A,x_{0},y_{0})$ correspond to source strength and location, which are the quantities of interest in practical applications such as seismic source localization Waldhauser and Ellsworth [2000] and electromagnetic inverse scattering Chen [2018].
To validate this theoretical analysis, we attempted direct field reconstruction on the full $64\times 64$ grid (4096 field values). We modified our framework so the inverse operator outputs 4096 values instead of 3 parameters, keeping all other architectural choices identical. We conducted two experiments: (1) MLP-based inverse operator: A multilayer perceptron maps sensor measurements directly to source field values. (2) DeepONet-based inverse operator: A DeepONet architecture where the branch network processes sensor locations and the trunk network processes evaluation points $(x,y)$.
Figures 27 and 28 present the results. For the MLP approach, the reconstructed source achieves relative $L_{2}$ error of $0.34$. While the inverse operator correctly identifies the approximate region of maximum source intensity, it fails to capture the detailed spatial structure of the true Gaussian source (Figure 27(a)). The solution field reconstructed from this source achieves relative $L_{2}$ error of $0.050$ (Figure 28(a)), indicating that multiple source configurations can produce similar wave fields at the sensor locations.
(a) MLP-based inverse operator: solution field
(b) DeepONet-based inverse operator: solution field
Figure 28: Helmholtz equation: Solution field comparison for non-parametric reconstruction. (a) Solution from MLP-reconstructed source with relative $L_{2}$ error of $0.0497$. (b) Solution from DeepONet-reconstructed source with relative $L_{2}$ error of $0.1523$. For each subfigure: left panel shows the reference FDM solution $u(x,y)$, middle panel shows the predicted solution, and right panel shows the pointwise absolute error. Multiple source configurations produce similar wave fields at sensor locations, demonstrating the non-uniqueness of the inverse problem.
The DeepONet-based inverse operator produces smoother reconstructions but exhibits spatial diffusion, with source error of $0.60$ (Figure 27(b)). The corresponding solution field error is $0.15$ (Figure 28(b)). The increased smoothness reflects the inherent bias of operator networks toward continuous functions but does not improve accuracy. Both architectures struggle to recover the true source structure despite different inductive biases.
These results confirm that the ill-posedness is a fundamental property of the Helmholtz inverse source problem, independent of the neural network architecture. Without explicit regularization through parametrization, smoothness penalties, or other priors, stable high-fidelity source reconstruction from sparse measurements is not achievable. The parametric approach remains the most appropriate choice for this problem class, combining mathematical well-posedness with physical interpretability.
3.5 Robustness of parameter identification
To evaluate the practical applicability of the proposed frameworks, we examine their robustness under two realistic scenarios: (i) noisy observations and (ii) sparse data availability. We analyze the parameter identification framework across representative benchmark problems.
Table 5 presents parameter identification accuracy under varying noise levels. Gaussian noise $\eta\sim\mathcal{N}(0,\sigma^{2})$ was added to sensor measurements during both training and testing, with $\sigma$ representing the percentage of signal standard deviation. The results demonstrate remarkable robustness enabled by physics-informed training. For 1D problems (Burgers’ and Reaction-Diffusion), parameter estimation maintains absolute errors on the order of $10^{-3}$ even at 5% noise, with relative solution errors below 12%. The Heat equation shows moderate degradation with thermal conductivity errors around $5.5\times 10^{-2}$ at 5% noise, while the Helmholtz source reconstruction achieves spatial localization errors below $5\times 10^{-2}$ across all noise levels, though amplitude estimation exhibits larger errors due to the parameter’s inherent ambiguity. Critically, solution reconstruction quality remains strong across all problems (relative $L_{2}$ errors <0.6<0.6 for Heat and <0.3<0.3 for Helmholtz), confirming that the physics-informed framework effectively regularizes the inverse problem through embedded governing equations, enabling stable parameter estimation from noisy sparse measurements.
Table 6 demonstrates the framework’s robustness to sparse sensor data. For Reaction-Diffusion and Burgers’ equations, both parameter identification accuracy and solution reconstruction improve monotonically with increasing sensor density, with relative $L_{2}$ errors decreasing from 0.088 to 0.050 and 0.048 to 0.029, respectively. The Heat equation exhibits strong robustness, with parameter error improving by over 50% (from $6.6\times 10^{-2}$ to $3.0\times 10^{-2}$) while maintaining consistent solution accuracy (relative $L_{2}$ error $\approx$ 0.57-0.71). The Helmholtz problem shows an interesting pattern: spatial parameters ($x_{0}$, $y_{0}$) remain accurately identified even at the sparsest configuration (64 sensors), while the amplitude parameter $A$ exhibits sensitivity to sensor density, with error increasing from $2.0\times 10^{-2}$ to $2.1$ at very sparse configurations. This reflects the higher sensitivity of wave field reconstruction to source amplitude under limited measurements. Importantly, even at the sparsest sensor configurations, all problems achieve parameter errors suitable for reasonable solution reconstruction, demonstrating the effectiveness of physics-informed training in data-limited scenarios.
Table 5: Parameter identification accuracy i.e. average absolute test error in the test cases under varying noise levels. Sensor measurements were corrupted with Gaussian noise $\eta\sim\mathcal{N}(0,\sigma^{2})$ where $\sigma$ represents the percentage of signal standard deviation. All experiments use 300 sensors.
Problem Parameter Noise Level (%)
0% 1% 3% 5%
Reaction-Diffusion $|D_{\text{true}}-D_{\text{pred}}|$ $8.0\times 10^{-3}$ $8.0\times 10^{-3}$ $9.0\times 10^{-3}$ $9.0\times 10^{-3}$
Rel. $L_{2}$ error ($u$) 0.050 0.057 0.111 0.112
Burgers’ $|\nu_{\text{true}}-\nu_{\text{pred}}|$ $2.0\times 10^{-3}$ $2.0\times 10^{-3}$ $3.0\times 10^{-3}$ $4.0\times 10^{-3}$
Rel. $L_{2}$ error ($u$) 0.029 0.037 0.044 0.049
Heat $|K_{\text{true}}-K_{\text{pred}}|$ $3.0\times 10^{-2}$ $5.6\times 10^{-2}$ $6.6\times 10^{-2}$ $5.5\times 10^{-2}$
Rel. $L_{2}$ error ($u$) 0.570 0.512 0.534 0.581
Helmholtz $|A_{\text{true}}-A_{\text{pred}}|$ $2.8\times 10^{-1}$ $2.0\times 10^{0}$ $2.5\times 10^{0}$ $2.1\times 10^{0}$
$|x_{0,\text{true}}-x_{0,\text{pred}}|$ $4\times 10^{-2} 3.3$$3\times 10^{-2}$ $2.6\times 10^{-2}$ $3.5\times 10^{-2}$
$|y_{0,\text{true}}-y_{0,\text{pred}}|$ $4.4\times 10^{-2}$ $3.7\times 10^{-2}$ $4.5\times 10^{-2}$ $3.9\times 10^{-2}$
Rel. $L_{2}$ error ($u$) 0.099 0.266 0.089 0.281
Table 6: Robustness of deterministic parameter identification to data sparsity. Results show mean absolute parameter error and relative $L_{2}$ error for the state variable $u$ over test samples under varying sensor counts. All measurements are noise-free.
Problem Metric Config 1 Config 2 Config 3 Config 4
Reaction-Diffusion $|D_{\text{true}}-D_{\text{pred}}|$ 50: $5.0\times 10^{-3}$ 100: $6.0\times 10^{-3}$ 200: $7\times 10^{-3}$ 300: $8.0\times 10^{-3}$
Rel. $L_{2}$ error ($u$) 0.088 0.077 0.073 0.050
[…]
Rel. $L_{2}$ error ($u$) 0.048 0.034 0.035 0.029
Heat $|K_{\text{true}}-K_{\text{pred}}|$ 165: $6.6\times 10^{-2}$ 231: $7.0\times 10^{-2}$ 352: $7.0\times 10^{-2}$ 440: $3.0\times 10^{-2}$
Rel. $L_{2}$ error ($u$) 0.71 0.59 0.56 0.570
Helmholtz $|A_{\text{true}}-A_{\text{pred}}|$ 64: $2.0\times 10^{-2}$ 144: $1.3\times 10^{-2}$ 256: $1.48\times 10^{-1}$ 1024: $2.1\times 10^{0}$
$|x_{0}^{\text{true}}-x_{0}^{\text{pred}}|$ $1.8\times 10^{-2}$ $1.8\times 10^{-2}$ $1.9\times 10^{-2}$ $3.5\times 10^{-2}$
$|y_{0}^{\text{true}}-y_{0}^{\text{pred}}|$ $2.9\times 10^{-2}$ $1\times 10^{-2} 2.2$$2\times 10^{-2}$ $3.9\times 10^{-2}$
Rel. $L_{2}$ error ($u$) 0.113 0.113 0.121 0.281
4 Discussions and Conclusion
4.1 Computational efficiency and scalability analysis
While the benchmark problems in this study are low-dimensional and individual forward simulations can be solved efficiently using traditional numerical methods (requiring $0.05$ to $0.12$ seconds depending on discretization), the proposed operator learning framework offers substantial computational advantages for inverse problems and parameter estimation tasks. Traditional approaches to parameter identification require iterative optimization, where each parameter estimation scenario involves multiple forward solver evaluations, typically $500$ or more iterations for convergence, along with gradient computation via finite differences or adjoint methods. For our benchmark problems, this translates to approximately $25$ to $60$ seconds per parameter estimation scenario, assuming a minimum of $500$ iterations. When parameter estimation must be performed across multiple scenarios $N$, the total computational cost scales as $\mathcal{O}(N\times n_{\text{iter}}\times t_{\text{forward}})$, making it prohibitively expensive for large-scale parameter space exploration. For instance, estimating parameters for $1,500$ test scenarios using traditional iterative optimization would require an estimated 10.410.4to $25$ hours of computation time, and this assumes relatively fast convergence without accounting for difficult optimization landscapes that may require significantly more iterations.
Table 7: Computational cost: training runtime and inference time for all models.
Problem DHPO INV-DON (Det.) INV-DON (Prob.) Inference
Train [hrs] Train [s] Train [s] [ms / query]
Reaction–Diffusion 1D 3 371.32 263.68 0.104
Burgers’ 1D 70 446.30 450.08 0.101
Heat 2D (L-shaped) 19 144.14 224.46 0.217
Helmholtz 2D (Source ID) 185.30 233.57 0.188
In contrast, our operator learning framework requires a one-time training phase ranging from $2.4$ to $7.5$ minutes depending on the problem complexity, after which parameter estimation for each new scenario requires only rapid inference (approximately $0.1$ to $0.24$ miliseconds per query) without any iterative optimization. As demonstrated in Table 7, it should be noted that DHPO training times are reported in hours, whereas INV-DON training times are reported in seconds, due to differences in implementation (TensorFlow 2 for DHPO versus JAX for INV-DON), hyperparameter choices (e.g., batch size, number of training samples, and epochs), and GPU hardware (V100 for DHPO versus A100 for INV-DON).These implementation differences do not affect the inference performance, which remains consistently fast across all methods. Our framework achieves parameter estimation for $1,500$ test scenarios in approximately $7.5$ minutes total time for the Burgers’ equation and $6.2$ minutes for the reaction diffusion system, representing speedups of over $80\times$ to $240\times$ compared to traditional iterative methods. For the 2D heat equation with $2,000$ test scenarios, the total time is approximately $3.7$ minutes compared to an estimated $13.9$ to $33.3$ hours using traditional optimization, yielding speedups of $225\times$ to $540\times$. This advantage becomes particularly pronounced for applications requiring real-time parameter estimation, extensive parameter space exploration for uncertainty quantification, or repeated inverse problems with different observation sets. We acknowledge that for single forward problem evaluations, traditional numerical methods are computationally efficient and well-established. However, the inverse problems addressed in this work, specifically parameter identification from sparse observations, require fundamentally different computational considerations, as traditional optimization based methods must repeatedly evaluate the forward solver whereas our operator learning framework learns the inverse mapping directly, enabling rapid parameter estimation across multiple scenarios without iterative optimization.
4.2 Conclusion
In this work, we presented two complementary operator-learning frameworks for solving inverse problems in computational science and engineering. The Deep Hidden Physics Operator (DHPO) extends hidden-physics modeling into the operator-learning paradigm, enabling the discovery of unknown PDE terms across diverse equation families. In parallel, the proposed parameter identification framework leverages pretrained DeepONets to reconstruct solution fields from sparse sensor data and infer governing parameters in a physics-consistent manner. Together, these approaches address two fundamental challenges, the hidden-physics discovery and parameter estimation within a unified data-efficient framework.
Our results on canonical benchmarks, including the reaction-diffusion system, Burgers’ equation, and the 2D heat equation, demonstrate that the methods achieve robust performance even under sparse and noisy observations. Specifically, DHPO successfully recovers hidden physics with relative solution errors on the order of $10^{-2}$, while the parameter identification framework achieves absolute parameter errors on the order of $10^{-3}$. Compared to existing methods such as PINNs and DHPM, the proposed frameworks offer improved generalization across PDE families, eliminate the need for repeated retraining, and enhance data efficiency.
These findings highlight the promise of combining neural operator learning with physics-informed constraints for advancing data-driven discovery in the sciences. Future directions include extending the frameworks to three-dimensional problems, incorporating uncertainty quantification for rigorous error bounds, and exploring hybrid architectures that integrate other neural operators (e.g., Fourier or wavelet-based operators) for enhanced scalability. More broadly, the operator-learning perspective opens new opportunities for robust, generalizable, and interpretable solvers in computational mechanics and beyond.
4.3 Limitations and future directions
While our approach demonstrates robust performance across various inverse problems, the current framework assumes fixed sensor locations during both training and inference. This represents a practical limitation, as many real-world inverse problems involve variable numbers and spatial distributions of sensors depending on measurement constraints or sensor availability. The fixed-location assumption is inherited from the DeepONet architecture, where the branch network processes sensor measurements at predetermined locations. Recent work on set-based operator networks (SetONet) Tretiakov et al. [2025] addresses this limitation by incorporating permutation-invariant set encoders that can handle arbitrary sensor configurations. Extending our Neural Inverse Operator framework to accommodate variable sensor placements represents a promising direction for future research, potentially combining our operator-to-function learning approach with SetONet’s flexible sensor encoding mechanisms. Such an extension would enhance the practical applicability of the method to scenarios with heterogeneous sensor networks or adaptive sensing strategies.
Acknowledgments
VK and BP would like to acknowledge the computing support provided by The Deep Learning and GPU computing Platform (BD/DPA-BDP1) at Robert Bosch, GmbH, Renningen, Germany. DRS and SG would like to acknowledge computing support provided by the Advanced Research Computing at Hopkins (ARCH) core facility at Johns Hopkins University and the Rockfish cluster. ARCH core facility (rockfish.jhu.edu) is supported by the National Science Foundation (NSF) grant number OAC1920103. DRS and SG would like to acknowledge Dr. Krishna Kumar, Dr. Xingjian Li, and Dr. Hassan Iqbal for the insightful discussions during the journal club meetings. DRS and SG were partially supported by JHU Discovery award 2024 and NSF grant number 2436738.
Author contributions
Conceptualization: VK, DRS, BP, SG
Investigation: VK, DRS, BP, SG
Visualization: VK, DRS, BP, SG
Supervision: BP, SG
Writing - original draft: VK, DRS
Writing - review & editing: VK, DRS, BP, SG
Data and code availability
The code and data that support the findings of this study will be available from the corresponding author on reasonable request upon publication.
Competing interests
The authors declare no competing interest
Appendix A Hyperparameter tuning
In the deep learning community, Weights & Biases (WandB) is a popular online platform that provides tools for monitoring and controlling hyperparameters during model training runs. Sweeps, a crucial component of WandB, allow for the systematic tuning of model hyperparameters. The Sweeps configuration dictionary contains the predefined hyperparameter search space. Depending on this setup, a sweep could include several runs, with grid search, random search, or Bayesian optimization used for hyperparameter sampling. Sarkar et al. Sarkar et al. [2025a] showcases the use of WandB for hyperparameter tuning on PINNs training. We have utilized the WandB tool for hyperparameter tuning of our framework for the parameter identification problem of the Burgers’ equation. The report of the hyperparameter tuning can be found at the link https://api.wandb.ai/links/droysar1-johns-hopkins-university/pzs4hfup.
Table 1: Architecture and training parameters for all the examples presented in this work.
Equation
Branch
Network
Trunk
Network
MLP or
FNN
Activation
Function
Training
Iterations
Learning Rate
(Initial / Decay)
Batch
Size
Hidden physics
Reaction-Diffusion $[101,128,128,128,50]$ $[2,128,128,128,50]$ $[3,128,128,128,1]$
Branch - ReLU
Trunk - Tanh
MLP - Tanh
10k 1e41e-4 10
Burgers’
$[101,128,128,$
$128,128,50]$
$[2,128,128,$
$128,128,50]$
$[3,256,256,256,1]$
Branch - ReLU
Trunk - Tanh
MLP - Tanh
10k 1e41e-4 1
Heat Equation (2D)
CNN: $33\times 33$,
Channels $[1,40,60]$,
Kernel $3\times 3$, stride 22,
Maxpool $2\times 2$
FCN: $[240,64,64,64]$
$[3,128,128,128,64]$ $[5,128,128,128,1]$
Branch - ReLU, Linear
Trunk - Tanh, Linear
MLP - Tanh, Linear
2000 1e41e-4 1
System parameter identification
Reaction-Diffusion $[300,64,64,64,100]$ $[2,64,64,64,100]$ $[300,64,64,64,1]$ Tanh 80k
1e31e-3
exponential,
rate = 0.9
$3500$
Burgers’ $[300,64,64,64,100]$ $[2,64,64,64,100]$ $[300,64,64,64,1]$ Tanh 80k
1e31e-3
exponential,
rate = 0.9
$2500$
Heat $[440,64,64,64,100]$ $[3,64,64,64,100]$ $[440,32,16,8,1]$
Branch-Tanh
Trunk-Tanh
MLP - GeLU
80k
1e31e-3
exponential,
rate = 0.9
$75$
Helmholtz $[1024,32,32,32,100]$ $[2,32,32,32,100]$ $[1024,32,32,32,3]$
Branch-Tanh
Trunk-Tanh
MLP - GeLU
80k
1e31e-3
exponential,
rate = 0.9
$75$
Figure 1: Loss function over the training process of optimal model: (a) Reaction diffusion equation with $N_{train}=500,N_{d}=500.$ (b) Burgers’ equation with $N_{train}=1000,N_{d}=500.$ (c) Heat equation with $N_{train}=500,N_{d}=550.$
Figure 2: Sample 1: hidden physics solution comparison, mean absolute error = 0.03177
Figure 3: Sample 2: hidden physics solution comparison, mean absolute error = 0.03415
(a) Sample 1: solution field accuracy comparison, mean absolute error $u_{pred}(t\leq 1)=0.00507,u_{pred}(t>1)=0.04674,\\ u_{hpm}(t\leq 1)=0.00627,u_{hpm}(t>1)=0.03299$
(b) Sample 2: solution field accuracy comparison, mean absolute error $u_{pred}(t\leq 1)=0.06620,u_{pred}(t>1)=0.13954,\\ u_{hpm}(t\leq 1)=0.08864,u_{hpm}(t>1)=0.35198$
Figure 4: Reaction diffusion equation: Comparison of DHPO solution & HPM integrated solution for two representative test samples with input function from GRF function space $l=0.2$ (nonzero at boundaries). It should be noted that training data is collected from t = 0 to 1., extrapolation time region separated by the white horizontal lines.
(a) Sample 1: solution field accuracy comparison, mean absolute error $u_{pred}(t\leq 1)=0.00458,u_{pred}(t>1)=0.05970,\\ u_{hpm}(t\leq 1)=0.0059,u_{hpm}(t>1)=0.03579$
(b) Sample 2: solution field accuracy comparison, mean absolute error $u_{pred}(t\leq 1)=0.02584,u_{pred}(t>1)=0.12762,\\ u_{hpm}(t\leq 1)=0.02967,u_{hpm}(t>1)=0.13450$
Figure 5: Reaction diffusion equation: Comparison of DHPO solution & HPM integrated solution for two representative test samples with input function from Modified GRF function space $l=0.2$ (zero at boundaries). It should be noted that training data is collected from t = 0 to 1., extrapolation time region separated by the white horizontal lines.
(a) Sample 1: hidden physics solution comparison, mean absolute error = 0.00567.
(b) Sample 2: hidden physics solution comparison, mean absolute error = 0.01787.
Figure 6: Reaction diffusion equation: Comparison of reference and predicted PDE terms $\mathcal{N}$ with input function from sine function space. It should be noted that, the Reference $\mathcal{N}$ is computed using finite difference with the gradients being computed using second order scheme.
(a) Sample 1: hidden physics solution comparison, mean absolute error = 0.01587.
(b) Sample 2: hidden physics solution comparison, mean absolute error = 0.21618.
Figure 7: Reaction diffusion equation: Comparison of reference and predicted PDE terms with input function from GRF function.
(a) Sample 1: hidden physics solution comparison, mean absolute error = 0.01435.
(b) Sample 2: hidden physics solution comparison, mean absolute error = 0.07512.
Figure 8: Reaction diffusion equation: Comparison of reference and predicted PDE terms with input function from Modified GRF function.
(a) Sample 1: hidden physics solution comparison, mean absolute error = 0.01267.
(b) Sample 2: hidden physics solution comparison, mean absolute error = 0.05602.
Figure 9: Burgers’ equation: Comparison of reference and predicted PDE terms.
Figure 10: DeepONet solution: Sparse training solution comparison. mean absolute error at training time instants $(t=0,1)$ is 0.03029, mean absolute error at testing time instants $(0<t<1)$ is 0.38671.